# Asymptotics of Proximity Operator for Squared Loss and Performance Prediction of Nonconvex Sparse Signal Recovery

### v3

@@ IV-D Performance Prediction of Douglas-Rachford Algorithm with Nonconvex Regularization
Figure 6: MSE performance of the Douglas-Rachford algorithm with nonconvex SCAD regularization ($\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\lambda=0.1$, $\gamma=1$, $\rho_{k}=1$, $a=4$).
The empirical performance and the prediction are obtained in the same way as Fig. 3. From Fig. 6, we observe that the empirical MSE and its prediction converge to almost the same value in this case. In the middle of iterations, however, the behavior of the prediction is different from Fig. 3, and the empirical performance when $N=500$ is better than the prediction. One of the possible reasons is the nonconvexity of the SCAD regularizer, and further investigation is necessary for a clear understanding.
Figure 7 shows the MSE at the $50$-th iteration versus the regularization parameter $\lambda$, where $N=500$, $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\gamma=1$, $\rho_{k}=1$, and $a=4$.
Figure 7: MSE performance of the Douglas-Rachford algorithm with nonconvex SCAD regularization versus $\lambda$ ($N=500$, $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\gamma=1$, $\rho_{k}=1$, $a=4$).
From the figure, we can see that the prediction is close to the empirical performance for $\lambda\geq 0.08$, whereas the prediction is worse for smaller values of $\lambda$. Even in this case, however, we can tune a reasonable value of $\lambda\approx 0.07$ on the basis of the prediction, without the empirical reconstruction.
@@ V Conclusion
In this paper, we have analyzed the asymptotic behavior of the proximity operator for the squared loss function in linear inverse problems. We have also shown that the result in Proposition III.1 can be used for the asymptotic performance prediction of some optimization algorithms for large-scale linear inverse problems. Simulation results show that the empirical distribution of the output of the proximity operator agrees well with the theoretical prediction by Proposition III.1. Moreover, the empirical MSE performance of the Douglas-Rachford algorithm can be well predicted in large-scale compressed sensing via the $\ell_{1}$ optimization. For the nonconvex SCAD regularization, the prediction is also valid to some extent, although the behavior is slightly different from the convex case.
Future work includes the rigorous analysis of the prediction by Conjecture III.1 and the further investigation of interesting behavior for the case with nonconvex regularization. Extensions of our approach beyond Assumptions III.1 and III.2 would also be an important research direction. Applications of our results to optimization algorithms other than the Douglas-Rachford algorithm would also be an interesting topic.

### v4

@@ IV-D Performance Prediction of Douglas-Rachford Algorithm with Nonconvex Regularization
Figure 9: MSE performance of Douglas-Rachford algorithm with nonconvex SCAD regularization ($\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\lambda=0.1$, $\gamma=1$, $\rho_{k}=1$, $a=4$).
The empirical performance and the prediction are obtained in the same way as Fig. 3. From Fig. 9, we observe that the empirical MSE and its prediction converge to almost the same value in this case. In the middle of iterations, however, the behavior of the prediction is different from Fig. 3, and the empirical performance when $N=500,1000$ is better than the prediction. One of the possible reasons is the nonconvexity of the SCAD regularizer, and further investigation is necessary for a clear understanding. To obtain more precise results, we might need to taking the nonconvexity into account in the analysis.
We also investigate the performance of the Douglas-Rachford algorithm with SCAD regularization for a larger noise variance $\sigma_{\text{v}}^{2}=0.01$ in Fig. 10, where $\Delta=0.7$, $p_{0}=0.9$, $\lambda=0.2$, $\gamma=1$, $\rho_{k}=1$, and $a=4$.
Figure 10: MSE performance of Douglas-Rachford algorithm with SCAD regularization ($\Delta=0.7,p_{0}=0.9,\sigma_{\text{v}}^{2}=0.01,\lambda=0.2,\gamma=1,\rho_{k}=1,a=4$).
From the figure, we can see that the empirical performance is well predicted in this case. The difference in behavior compared to Fig. 9 is probably influenced by the value of $\sigma_{\text{v}}^{2}$ and other parameters, presenting an intriguing avenue for future research.
Figure 11 shows the performance of the Douglas-Rachford algorithm with the nonconvex MCP regularization.
Figure 11: MSE performance of Douglas-Rachford algorithm with nonconvex MCP regularization ($\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\lambda=0.1$, $\gamma=1$, $\rho_{k}=1$, $b=4$).
In the simulation, we set $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\lambda=0.1$, $\gamma=1$, $\rho_{k}=1$, and $b=4$. As in Fig. 9, the empirical performance when $N=500,1000$ is better than the prediction in the middle of iterations. These results suggest that the nonconvexity of the MCP regularizer also affects the behavior of the prediction.
Figure 12 shows the MSE at the $50$-th iteration versus the regularization parameter $\lambda$, where $\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\gamma=1$, and $\rho_{k}=1$.
Figure 12: MSE performance of Douglas-Rachford algorithm with nonconvex SCAD regularization versus $\lambda$ ($\Delta=0.7$, $p_{0}=0.9$, $\sigma_{\text{v}}^{2}=0.001$, $\gamma=1$, $\rho_{k}=1$, $a=4$, $k=50$).
In the simulation, we use the SCAD regularization with $a=4$. From the figure, we can see that the prediction is close to the empirical performance for $\lambda\geq 0.08$, whereas the prediction is worse for smaller values of $\lambda$. Even in this case, however, we can tune a reasonable value of $\lambda\approx 0.07$ on the basis of the prediction, without the empirical reconstruction.
@@ V Conclusion
In this paper, we have analyzed the asymptotic behavior of the proximity operator for the squared loss function in linear inverse problems. We have also shown that the result in Proposition III.1 can be used for the asymptotic performance prediction of some optimization algorithms for large-scale linear inverse problems. Simulation results show that the empirical distribution of the output of the proximity operator agrees well with the theoretical prediction by Proposition III.1. Moreover, the empirical MSE performance of the Douglas-Rachford algorithm can be well predicted in large-scale compressed sensing via the $\ell_{1}$ optimization. For the nonconvex SCAD regularization and MCP regularization, the prediction is also valid to some extent, although the behavior is slightly different from the convex case. These results suggest that the proposed prediction method can be a foundation of the rigorous analysis of optimization algorithms for large-scale linear inverse problems.
Future work includes the rigorous analysis of the prediction by Conjecture III.1 and the further investigation of interesting behavior for the case with nonconvex regularization. One of possible directions is to extend or modify the approach in [40], though independent measurement matrices at each iteration of the algorithm is assumed in their analysis. Extensions of our approach beyond Assumptions III.1 and III.2 would also be an important research direction. Applications of our results to optimization algorithms other than the Douglas-Rachford algorithm would also be an interesting topic.